x+1 4x – 7 c) f(x)= ; c= 0 2x2 + 3x – 2 d) f(x)= 2x2 + 5x – 3 ; C = 2

Find a power series for the function, centered at c, and determine the interval of convergence.
Answer #1

We first express both the function into standard form for the power series expansion and then we find the interval of convergence.

It is to be noted that both the power series expansion constant are in the summation form not in product form so to perform ratio (or root) test we separately calculated the intervals and finally we took the intersection of the interval for interval of convergence.

like ratio test) Now we calculate the limit (almost (-1) + 3 2 n+1 ( First passt of series), oh lim no I N12 (-1) n. 3 m 2 12f(x) fly) W 90 @ family X+1 ; C = 2x² +5 x-3 2x24 5x-3 f(x) 2x² + 6x-2x-3 (x-1) (x+3) = (2x+1)(2+3) Now X+1 O Powen R+1) 13/12 35 t PO 1+ 2y It 3 5 n=0 i fly) = lomoto power Semes will be sobe)= 2 (o6333 +ão conclusion [***. 4 + şi loom] yn poo [6